OpenStudy (anonymous):

PLEASE HELP ME AND I WILL GIVE A MEDAL! Refer to the following conditional statement: If a and b are odd integers, then a-b is an even integer. Identify the conclusion. 1. a and b are not odd integers 2. a and b are odd integers 3. a-b is an even integer. 4. a-b is not an even integer

OpenStudy (anonymous):

@Here_to_Help15

OpenStudy (anonymous):

@lacrosseplayer22

OpenStudy (anonymous):

@BassCatcher15

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

Pls help

OpenStudy (princeharryyy):

2 and 3 are correct always true

OpenStudy (anonymous):

But which one is it? I can only pick one

OpenStudy (princeharryyy):

then choose 2nd

OpenStudy (princeharryyy):

a-b is even only if a & b are both odd

OpenStudy (anonymous):

Refer to the following conditional statement:

OpenStudy (anonymous):

Okay thank you! I have some more if you could help me please?

OpenStudy (princeharryyy):

example ; 7-3 = 4 ans , ur welcome ask ?

OpenStudy (princeharryyy):

be quick i've other stuff to do ...

OpenStudy (anonymous):

Refer to the following statement: if a and b are odd integers, then a * b is not an odd integer. State the converse of the statement. 1. If a and b are not odd integers then a * b is not an odd integer. 2.if a * b is an odd integer then a and b are odd integers 3. a and b are odd integers 4. a * b is an odd integer

OpenStudy (anonymous):

Sorry that took so long I'm typing this on my iPad :( lol

OpenStudy (anonymous):

5. Refer to the following conditional statement: if a and b are odd integers then a- b is an even integer. Sate the inverse of the statement. 1. If a-b is an even integer then a and b are odd integers 2. a-b is an even integer 3. if a-b is not an even integer then a and b are not odd integers 4. If a and b are not odd integers then a-b is not an even integer

OpenStudy (anonymous):

Are you still there?

OpenStudy (anonymous):

@aum

OpenStudy (aum):

if a-b is not an even integer then a and b are not odd integers

OpenStudy (anonymous):

Thank you!! Do you think you could also answer the one before that please?

OpenStudy (anonymous):

@aum

OpenStudy (princeharryyy):

hmmm wait was doing some work of my own...

OpenStudy (princeharryyy):

let me check ur questions

OpenStudy (anonymous):

Okay thank you :)

OpenStudy (princeharryyy):

Refer to the following statement: if a and b are odd integers, then a * b is not an odd integer. State the converse of the statement. 1. If a and b are not odd integers then a * b is not an odd integer. 2.if a * b is an odd integer then a and b are odd integers 3. a and b are odd integers 4. a * b is an odd integer ans>> if a*b is not an odd integer than a and b are odd integers

OpenStudy (princeharryyy):

are you sure that the question statement that u wrote is correct because that statement is wrong it can never happen

OpenStudy (anonymous):

I'm not sure let me check

OpenStudy (princeharryyy):

because multiplication of two odd integers always yields an odd integer as a result

OpenStudy (anonymous):

Oh we'll I'm not sure I checked and those are the options they gave me

OpenStudy (princeharryyy):

its not about the options its about concept see 1*3 =3, 3*5=15, 5*7=35, 7*9=63 and so on, see result is always an odd number... its universal

OpenStudy (princeharryyy):

hey hello r u there ????

OpenStudy (anonymous):

Sorry I had to make food for my brother

OpenStudy (anonymous):

Okay so what would the answer be then?

OpenStudy (princeharryyy):

do you know wat converse mean in mathematics

OpenStudy (anonymous):

Nope

OpenStudy (princeharryyy):

ohk le' me explain u with an example

OpenStudy (princeharryyy):

suppose that i say a+b=5, then if i say, if a,b are 2,3 or 3,2 then the sum is always 5.

OpenStudy (princeharryyy):

its converse would be if the sum is 5 than a,b should be 2,3 or 3,2

OpenStudy (princeharryyy):

so converse means if the first statement complies with the second then the second would also comply with first one

OpenStudy (anonymous):

Oh okay!

OpenStudy (princeharryyy):

now, would you be able to do the similar questions

OpenStudy (princeharryyy):

???? r u dere ?

OpenStudy (anonymous):

I don't think so lol and yes I am

OpenStudy (princeharryyy):

hehe )))and for the inverse ... seee here's an example >> Statement: If it rains, the game will be canceled. Inverse (Negate both parts): If it doesn't rain, the game will not canceled.

OpenStudy (anonymous):

Oh okay that makes sense but what about the other question. I thought about what you said about converse and I don't think there is a correct answer for it or option I mean lol

OpenStudy (princeharryyy):

well the question itself is wrong i can't do anything about it i can just tell u the concept i wrote u the write statement with an example accordingly u can make the answer

OpenStudy (princeharryyy):

???/

OpenStudy (anonymous):

Okay

OpenStudy (princeharryyy):

can i ask u something

OpenStudy (anonymous):

It would be number 2! I think . And sure lol

OpenStudy (princeharryyy):

hey where are u from

OpenStudy (anonymous):

USA lol wa

OpenStudy (anonymous):

Why? And would number 2 be right?

OpenStudy (princeharryyy):

ohk

OpenStudy (princeharryyy):

if a and b are odd integers, then a * b is not an odd integer => if a*b is not an odd integer then a and b are odd integers, match this with your options

OpenStudy (princeharryyy):

nothing as such with your place ))

OpenStudy (anonymous):

Okay thanks

OpenStudy (anonymous):

Where are you from?

OpenStudy (princeharryyy):

india well i don't use this forum if u need any help msg me at www.facebook.com/princeharryyy

OpenStudy (princeharryyy):

it may be possible i won't be using it again ))))